### v2

# bibliography.bib

Abstract
A common step at the core of many RNA transcript assembly tools is to find a set of weighted paths that best explain the weights of a DAG. While such problems easily become NP-hard, scalable solvers exist only for a basic error-free version of this problem, namely minimally decomposing a network flow into weighted paths.
The main result of this paper is to show that we can achieve speedups of two orders of magnitude also for path-finding problems in the realistic setting (i.e., the weights do not induce a flow). We obtain these by employing the safety information that is encoded in the graph structure inside Integer Linear Programming (ILP) solvers for these problems. We first characterize the paths that appear in all path covers of the DAG, generalizing a graph reduction commonly used in the error-free setting (e.g. by Kloster et al. [ALENEX 2018]). Secondly, following the work of Ma, Zheng and Kingsford [RECOMB 2021], we characterize the sequences of arcs that appear in all path covers of the DAG.
We experiment with a path-finding ILP model (least squares) and with a more recent and accurate one. We use a variety of datasets originally created by Shao and Kingsford [TCBB, 2017], as well as graphs built from sequencing reads by the state-of-the-art tool for long-read transcript discovery, IsoQuant [Prjibelski et al., Nat. Biotechnology 2023]. The ILPs armed with safe paths or sequences exhibit significant speed-ups over the original ones. On graphs with a large width, average speed-ups are in the range 50160×50-160$\times$50 - 160 × in the latter ILP model and in the range 1001000×100-1000$\times$100 - 1000 × in the least squares model.
Our scaling techniques apply to any ILP whose solution paths are a path cover of the arcs of the DAG. As such, they can become a scalable building block of practical RNA transcript assembly tools, avoiding heuristic trade-offs currently needed on complex graphs.
Keywords:
Path cover Network flow Flow decomposition Integer Linear Programming RNA-seq Genome-guided transcript assembly
1 Introduction
Background and motivation.
The genome-guided RNA transcript assembly problem, one the most famous assembly problems in bioinformatics, can be succinctly described as follows. Given a set of RNA-seq reads, a directed acyclic graph (DAG) is constructed from their alignments to a reference genome. The graph nodes correspond to e.g. exons, the arcs correspond to reads overlapping two consecutive exons, and the node or arc weights corresponding their read coverage. The RNA transcripts then correspond to a set of source-to-sink weighted paths in the DAG that “best explain” the nodes, arcs and their weights [tomescu2015explaining], under various definitions of optimality.
On perfect, error-free data, the arc weights satisfy flow conservation. In this setting, the most well-known definition of optimality is to require a minimum number of weighted paths whose superposition fully equals to the given flow weights (minimum flow decomposition, or MFD). This is a classical NP-hard problem [ahuja1988network, vatinlen2008simple], with many applications also in other fields, such as transportation [olsen2022study] and networking [hartman2012split, vatinlen2008simple]. On real data, the arc weights do not satisfy flow conservation because of errors in the reads and in their alignment, biases in sequencing coverage, and trimming in the reads, see e.g. [lrgasp]. As such, many practical tools model the RNA transcript assembly problem as a path-finding Integer Linear Program (ILP), for which they use a fast solver, such as Gurobi [gurobi] or CPLEX [cplex2009v12]. ILP is a powerful paradigm to model and efficiently solve NP-hard problems, including in bioinformatics, see e.g. [gusfield2019integer]. RNA transcript assembly based on ILP include CIDANE [cidane], CLASS2 [class2], TransLiG [translig], CLIIQ [cliiq], IsoInfer [isoinfer], IsoLasso [isolasso], MultiTrans [multitrans], NSMAP [nsmap], SSP [ssp], JUMPER [jumper].
Most of the existing ILP-based RNA transcript assembly tools (e.g. [cidane, class2, cliiq, isoinfer, isolasso, multitrans, nsmap, ssp, translig]) in principle do not scale with large graphs. The reason is that one first needs to enumerate all possible paths in the graphs, and then add an ILP variable for each path. This possibly leads to an exponential pre-processing time, and to an exponentially-sized ILP. Thus, many of these tools e.g. [cidane, cliiq, ireckon, isoinfer, multitrans, nsmap] use the heuristic of enumerating only some of all possible paths, potentially leaving some transcripts undiscovered, or leading to incorrect answers. Recently, [dias2022fast, jumper] observed that the enumeration step can be avoided by modeling the search for paths in the ILP itself, via only polynomially-many additional variables and constraints. However, some datasets still require tens of hours to solve [acceleratingILP, dias2024robust].
Despite this pressing need for fast solutions to path-finding ILPs modeling real-world data, most research effort has been put in the error-free setting, namely in the MFD problem. This includes fast heuristics [shao2017theory, vatinlen2008simple, hartman2012split, bernard2014efficient], fixed-parameter tractable algorithms [kloster2018practical], and approximation algorithms [hartman2012split, caceres2024width]. Recently, [acceleratingILP] showed that also the above-mentioned polynomially-sized ILP models for the MFD problem can be sped-up using some insight into the input flow structure, via the notion of safety [omnitigs_tomescu] (which we also review below). These optimizations also apply to an MFD variant where we are also given subpath constraints corresponding to long-reads aligned to the graph. On the hardest instances this leads to speedups of two orders of magnitude [acceleratingILP].
Contributions.
In this paper we show that similar speedups can be obtained also for real-world inputs where the arc weights do not satisfy flow conservation, and for any “path-finding” problem formulation, as long as their solution paths are a path cover of the arcs of the DAG. [^1] In this paper we assume that the input DAG has a unique source node $s$ and a unique sink node $t$; if this is not the case, one can just add a new global source $s$ connected to all existing graph sources (and symmetrically for a global sink), and specially handle these extra arcs in the problem formulations. Moreover, by a path cover we mean a set of paths from $s$ to $t$ ($s$-$t$ paths), such that every arc belongs to at least one path. We obtain these speedups by exploiting the graph structure of the DAG, in particular by exploiting the safe paths (and safe sequences of arcs) that appear in all path covers of the DAG. As such, we prove new results about the structure of directed acyclic graphs, which may also be of independent interest. More specifically, we give the following contributions.
1.0.1 1. Generalizing the Y-to-V reduction as finding safe paths for path covers.
For the minimum flow decomposition problem, Kloster et al. [kloster2018practical] used a graph reduction operation that decreases the size of the DAG, while preserving all flow decompositions. Namely, as long as the graph has a node $v$ with only one in-neighbor $u$, one removes $v$, and adds arcs from $u$ to each out-neighbor $w$ of $v$, with flow value $f(u,w)=f(v,w)$. This is a correct operation for MFD because of flow conservation: intuitively, the flow on any arc $(v,w)$ must come to $v$ via its unique in-neighbor $u$. A symmetric operation applies to nodes with only one out-neighbor (see Figure 4 (a) in Appendix 0.B). These operations have later been used in other works on the MFD problem, see [acceleratingILP, dias2023safety]. Moreover, they are also common in the context of the genome assembly problem [medvedev2007computability, jackson2009parallel, kingsford2010assembly], where this operation was called the “Y-to-V reduction” [omnitigs_tomescu] because of the shape of the subgraphs before and after reduction. We will also use this name to refer to this operation in this paper.
(a) Four safe paths, shown as colored lines, such that there is no $s$-$t$ path containing any two of them.
(b) Four safe sequences such that there is no $s$-$t$ path containing any two of them. Dotted lines indicate gaps in the sequences.
Figure 1: Example of safe paths and safe sequences for the $s$-$t$ path covers in a DAG $G$ with unique source $s$ and unique sink $t$. That is, for every path cover of $G$ (i.e. set of paths from $s$ to $t$ in $G$ such that every arc of $G$ appears in a path of the cover), there is a path in the cover containing the safe path or the safe sequence. We can see that safe sequences extend safe paths over complex subgraphs. If we have a set of safe path, or safe sequences, respectively, such that no two of them can appear on the same path of the DAG, then each of them must clearly appear in different paths of any path cover. Suppose that an ILP model has $k$ binary variables $x_{uvi}$, $i\in\{1,\dots,k\}$, for every arc $(u,v)$ of $G$, with the interpretation that $(u,v)$ appears in solution path $i$ iff $x_{uvi}=1$. Then, using such set of safe paths or sequences we can fix to 1 some of these binary variables, as in [acceleratingILP]. For example, we can assign the blue, orange, green and violet safe paths to solution paths 1,2,3 and 4, respectively (we show this assignment only for the blue and violet safe paths). We can proceed in a similar manner for safe sequences, now fixing more variables because sequences are overall longer (we show this assignment only for the blue and violet safe sequences)
However, when flow conservation does not hold, and when considering other path-finding problems than MFD, these are invalid operations. As such, we generalize the Y-to-V reduction so that we can use it for any problem whose solution is some path cover of the arcs. Specifically, we say that a path $P$ is safe (with respect to the path covers of a DAG $G$) if for any path cover of $G$, there is a path in the path cover that contains $P$ as subpath (see Figure 1(a) for an example). Using this framework, one can see the correctness of the Y-to-V reduction because the paths $(u,v,w)$, for any out-neighbor $w$ of $v$ are safe. However, by repeatedly applying the Y-to-V reduction, one misses safe paths (see Figure 4), and thus this reduction is not a correct algorithm for finding all safe paths. In this paper, we characterize all the safe paths with respect to the path covers of a DAG $G$ (Section 2). Using this characterization, we show that all maximal safe paths (i.e., those that contain all other safe paths as subpaths) can be computed in optimal linear time:
{restatable*}
[Maximal safe paths enumeration]theoremmaxsafepathsenum Given a DAG $G$ with $m$ arcs, there is an $O(m+o)$-time algorithm computing all the maximal safe paths of $G$, where $o$ is the total length of the output, namely of all maximal safe paths.
1.0.2 2. Generalizing safe paths to safe sequences for path covers.
While safe paths fully capture contiguous safety information, it may be that the only way to reach a safe path $P_{2}$ is by passing through a safe path $P_{1}$, with some “complex” subgraph between $P_{1}$ and $P_{2}$. As such, the sequence $P_{1},P_{2}$ appears in some path of any path cover. Therefore, inspired by the results of Ma, Zheng and Kingsford [and_or_quant], we generalize the previous notion of safe paths to safe sequences of arcs that appears in all paths covers of a DAG $G$ (see Figure 1(b) for an example). More specifically, an application of the AND-Quant problem from [and_or_quant] is to characterize when a sequence $(e_{1},\dots,e_{t})$ of arcs of $G$ appears in all flow decompositions of $G$. Namely, when it holds that for any flow decomposition, there is a path $P_{i}$ of the decomposition such that $e_{1},\dots,e_{t}$ appear in $P_{i}$, in this order. This is based on solving a max flow problem from $(e_{1},\dots,e_{t})$ and $G$. In this paper we characterize safe sequences with respect to path covers of $G$. For this, we generalize our results for safe paths by using the notion of $u$-$v$ bridges.
{restatable*}
[Maximal safe sequences enumeration]theoremmaxsafeseqsenum Given a DAG $G$ with $m$ arcs, there is an $O(m+o^{3})$-time algorithm computing a set of safe sequences containing all the maximal safe sequences of $G$, where $o$ is the total length of the output, namely of all maximal safe sequences.
1.0.3 3. Scaling path-finding ILPs via safe paths and sequences.
Safety characterizes the information that must appear in all solutions to a problem. If the solutions to a real-world path-finding problem are a subset of all the path covers of a DAG, then safe paths and safe sequences for path covers also appear in the solution paths of our path-finding problem. As such, we can use them to simplify a solver for the problem. For this, we use the approach of Grigorjew et al. [acceleratingILP]: if we are given a set of paths (or sequences, in our case) that must each be used by different solution paths, then we can fix some binary variables in the ILP model that encode the solution paths. See Figure 1 for an illustration.
We apply safe paths and safe sequences in this manner for two path-finding ILPs. In the first one (LeastSquares), we need to find a set of paths minimizing the sum of the squared errors between the weight of each arc and the weight of the solutions paths going through the arc. This is at the core of several RNA assembly tools e.g. [isolasso, cidane, ryuto, slide, ireckon, traph]. The second one (MinPathError) was recently introduced in [dias2024robust] and shown to be more accurate than LeastSquares (and than other ones, such as minimum inexact flow decomposition [inexact], see [dias2024robust]). The goal here is to account for errors not at the level of individual arcs, but at the level of solution paths, and minimize the sum of the errors of the paths. We describe all the above in detail in Appendix 0.B.
The ILPs optimized with safe paths or sequences exhibit significant speed-ups over the original ones, with average speed-up of at least 10×10$\times$10 × on many types of graphs. On graphs with a large width, average speed-ups are in the range $50\times$ 150 × for MinPathError, and in the range 1001000×100-1000$\times$100 - 1000 × for LeastSquares. As such, our optimizations can become a scalable building block of practical RNA transcript assembly tools, avoiding heuristic trade-offs on complex graphs currently needed.
2 Safe paths for path covers in $s$-$t$ DAGs
In this paper we work with DAGs with a unique source $s$ and a unique sink $t$, without loss of generality. We call such a graph an $s$-$t$ DAG. For an $s$-$t$ DAG $G=(V,E)$ we let $n=|V|$ and $m=|E|$. We adopt the following conventions: letters $u,v,x,y,z,w$ denote nodes, letters $a,b,h$ denote arcs, letters $p,q,r$ denote paths, and capital letters $A,B$ denote sequences of arcs.
Definitions.
We denote a path as a sequence of nodes $p=v_{1}v_{2}\dots v_{k}$. We say that a path $p$ covers (or contains) an arc $a$ if $a=v_{i}v_{i+1}$ for some $i\in\{1,\dots,k-1\}$, and we write $a\in p$ for that matter. We also write $v_{i}\in p$ if $v_{i}$ is in the node sequence of $p$. We let $V(p)$ and $E(p)$ denote the node and arc sets of $p$. We define an $s$-$t$ path cover of $G$ to be a set $P$ of $s$-$t$ paths such that $\forall a\in E~{}\exists p\in P:a\in p$. We say that a path $p$ is safe for $G$ if for any $s$-$t$ path cover $P$ of $G$ it holds that $p$ is a subpath of some path in $P$. The set of inner nodes of $p$ is the set $\{v_{2},\dots,v_{k-1}\}$ (if $p$ has 2 nodes or less then it contains no inner nodes). The indegree of a node $v$ is denoted by $d^{-}_{v}$ and the outdegree by $d^{+}_{v}$. If $q=u_{1}\dots u_{\ell}$ is a path where $v_{k}=u_{1}$, then $pq:=v_{1}\dots v_{k-1}u_{1}\dots u_{\ell}$ denotes their concatenation. For $1\leq i\leq j\leq k$, we write $\bar{v_{i}}p:=v_{i}\dots v_{k}$, $p\bar{v_{j}}:=v_{1}\dots v_{j}$, and $\bar{v_{i}}p\bar{v_{j}}:=v_{i}\dots v_{j}$. We say that node $v_{i}$ precedes (succeeds) node $v_{j}$ in $p$ if $i<j$ ($i>j$). We say that an arc $v_{i}v_{i+1}$ precedes (succeeds) an arc $v_{j}v_{j+1}$ in $p$ if $i<j$ ($i>j$), for $i,j\in\{1,\dots,k-1\}$. A path $p$ is maximally safe if it is not a subpath of any other safe path.
In this section we will show how to solve the following problem.
Problem 1
Given an $s$-$t$ DAG, find all its maximal safe paths.
Characterisation of safe paths.
The key observation is that a path $p=v_{1}\dots v_{k}$ containing an inner node $v_{i}$ of indegree at least two not succeeding an inner node $v_{j}$ of outdegree at least two is unsafe, as any path covering $p$ completely can be broken down into two $s$-$t$ paths: one covering the subpath $v_{1}\dots v_{j}$ without traversing $v_{j+1}\dots v_{k}$, and the other covering the subpath $v_{i}\dots v_{k}$ without traversing $v_{1}\dots v_{i-1}$. These two paths cover at least as many nodes as the original path containing $p$ and none contain $p$ as a subpath. The following theorem formalizes this intuition; the proof can be found in Appendix 0.A.
{restatable*}
theoremsafepathscharacterisation A path $p$ is safe if and only if any inner node $u$ with $d^{+}_{u}\geq 2$ precedes any inner node $v$ with $d^{-}_{v}\geq 2$ in $p$.
Having proven Section 2, we know that any safe path contains an arc separating the nodes with outdegree at least two from the nodes with indegree at least two.
Corollary 1
Any safe path $p$ contains an arc $h$ such that any inner node $u$ with $d^{+}_{u}\geq 2$ precedes $\textsc{h}(h)$ and any inner node $v$ with $d^{-}_{v}\geq 2$ $p$ succeeds $\textsc{t}(h)$ in $p$.
Computing safe paths via univocal extensions.
In order to compute every safe path we use the notion of univocal extension, first proposed in [schmidt2023omnitig], for which we give a functional definition. Given a safe path $p=v_{i}v_{i+1}\dots v_{j-1}v_{j}$, its right univocal extension is the right univocal extension of the concatenation of $p$ with the unique out-neighbour of $v_{j}$ (if $v_{j}$ has exactly one out-neighbor) or simply $p$ otherwise (thus stopping the recursion). Analogously, the left univocal extension is defined with respect to the in-neighbourhood of $v_{i}$. Note that the univocal extensions of a safe path (namely, an arc) are unique and never extend to the right over a node with outdegree at least two and never extend to the left over a node with indegree at least 2, and so by Section 2 the resulting path is safe. Moreover, by Corollary 1, any safe path can be seen as the univocal extension of some arc in the path or as a subpath of the extension [^2] For example, a proper subpath of an $s$-$t$ unitig path cannot be identified by univocal extensions..
Corollary 2
There are at most $m$ maximal safe paths.
The notion of univocal extension is computational in nature and directly gives us an algorithm to find safe paths: for every arc in the graph, compute its univocal extensions. Some resulting safe paths may not be maximal, see for example Figure 4 (a) and compare the extension of the arc $(u,v)$ with the extension of the arc $(v,w_{1})$; also, some safe paths could be repeated, for example, two arcs in a unitig have the same univocal extension. Nonetheless, since every safe path is the univocal extension of some arc, we are guaranteed to find all the maximal safe paths in the process. The univocal extension of an arc can be performed in $O(n)$ time, and so this algorithm runs in $O(mn)$ time.
To avoid computing duplicated and non maximal safe paths we use the notion of core from [schmidt2023omnitig]. A core is a unitig such that its univocal extension results in a maximal safe path. With this, we proceed as follows. We contract all unitigs into an arc to obtain a contracted graph; clearly, there is a bijection between safe paths on the contracted graph and safe paths on the original graph. To compute the set of maximal safe paths we perform univocal extensions in every arc of the contracted graph whose tail has outdegree at least two (or is the source) and whose head has indegree at least two (or is the sink), while keeping a left and right pointer in the original graph that mimic the extensions on the contracted graph: initially, the left pointer points to the first node of the core and the right pointer points to its last node, and for every arc $a$ that we extend to in the contracted graph (to the left or right) the corresponding pointer extends over the whole unitig in the original graph corresponding to $a$. We refer to [schmidt2023omnitig] for the proofs of correctness of this procedure. The contraction step can be done in $O(m+n)$ time and the univocal extension phase can be done in $O(m+o)$ time where $o$ is the total length of the maximal safe paths, all requiring only linear space in the size of the graph. This algorithm solves Problem 1 in optimal time and space.
[…]
3 Safe sequences for path covers in $s$-$t$ DAGs
We now generalise the previous results to safe sequences. The motivation behind this idea comes from the fact that some arcs must be in the same path of any path cover even though they may not appear consecutively in the path. For example, a set of non consecutive $s$-$t$ bridges of an $s$-$t$ DAG $G$ forms a safe sequence in our problem since any $s$-$t$ path must contain all the bridges of $G$. Note that we only impose that some path in the cover contains a given sequence of arcs, which is a weaker requirement.
Definitions.
Let $G=(V,E)$ be an $s$-$t$ DAG. For $A\subseteq E$ we write $A=a_{1},\dots,a_{\ell}$ as a sequence of arcs if there is an $a_{i}$-$a_{i+1}$ path for all $11\leq i\leq\ell-$ 1. We say that $A$ covers (or contains) an arc $b$ if $b=a_{i}$ for some $i\in\{1,\dots,\ell\}$, and we write $b\in A$ for that matter. If $B$ is a sequence and every arc of $A$ is contained in $B$ we say that $A$ is a subsequence of $B$ (since $G$ is a DAG, the arcs of $A$ appear in the same order as in $B$). Since a path $p=v_{1}\dots v_{k}$ naturally defines a sequence of arcs, we say that $A$ is a subsequence of $p$ or that $p$ covers $A$ if $A$ is a subsequence of $(v_{1},v_{2}),(v_{2},v_{3}),\dots,(v_{k-1},v_{k})$. We say that $A$ is safe for $G$ if in any $s$-$t$ path cover $P$ of $G$ there is a path $p\in P$ such that $A$ is a subsequence of $p$. A sequence $A$ is maximally safe if it is not a subsequence of any other safe sequence.
We let $R^{+}(u):=\{v\in V:\exists\text{$u$-$v$ path}\}$ and $R^{-}(u):=\{v\in V:\exists\text{$v$-$u$ path}\}$. For arcs $a=(x,y)$ and $b=(z,w)$, we let $R_{ab}:=R^{+}(y)\cap R^{-}(z)$. We say that a node $u$ is a fork (merge) for $ab$ if it is in $R_{ab}$ and there is a node $v\notin R_{ab}$ reachable from $u$ (reachable from $v$). We also let $R_{A}:=\bigcup_{i=1}^{\ell-1}R_{a_{i}a_{i+1}}$, and so the set of forks (merges) of $A$ becomes the union of all the forks (merges) in $R_{a_{i}a_{i+1}}$ for all $i\in\{1,\dots,\ell-1\}$. Observe that if $u,v\in R_{A}$ and $u$ reaches $v$, then, with respect to the sequence $A$, if $u$ is a merge then $v$ is a merge and if $v$ is a fork then $u$ is a fork. Without loss of generality, we assume that there is always an $s$-$t$ path containing a given sequence of arcs $A$, since otherwise $A$ is vacuously unsafe. Recall that for nodes $u,v\in V$, a $u$-$v$ bridge is an arc $b$ such that there is no $u$-$v$ path in $G-b$; equivalently, all $u$-$v$ paths in $G$ contain $b$.
Our goal now is to solve the following problem.
Problem 2
Given an $s$-$t$ DAG, find all its maximal safe sequences.
Characterisation of safe sequences.
The idea of our characterisation for safe sequences is analogous to that of safe paths, but where we lift the space to perform splices from the internal nodes of a path to $R_{A}$. See Appendix 0.A for a proof of the next theorem.
{restatable*}
theoremsafesequencescharacterisation A sequence of arcs $A=a_{1},\dots,a_{\ell}$ is safe if and only if there is a $\textsc{h}(a_{1})$-$\textsc{t}(a_{\ell})$ path $p$ containing all arcs in $a_{2},\dots,a_{\ell-1}$ such that any fork of $A$ in $p$ precedes any merge of $A$ in $p$.
Computing safe sequences via bridges.
Let $h=(u,v)$ be an arc, $A$ the sequence of $s$-$u$ bridges, and $B$ the sequence of $v$-$t$ bridges. We argue that the concatenation of these sequences ordered by reachability is a safe sequence. First, it is a sequence, as the last arc of $A$ (if any) reaches $h$ and $h$ reaches the first arc of $B$ (if any). Secondly, some $s$-$t$ path $q$ of any path cover must cover $h$, and since, by construction, any $v$-$t$ path contains $B$ and any $s$-$u$ path contains $A$, $q$ necessarily contains the sequence $AhB$. This shows that bridge extensions are safe.
Furthermore, as a consequence of Section 3, if $A$ is a safe sequence then there is special path $p$ containing an arc $h=(u,v)$ separating the forks from the merges of $A$ in $p$. Now, the sequence of $s$-$u$ bridges must contain the arcs of $A$ preceding $h$ in $p$, since otherwise there would be a merge in the nodes of $p$ preceding $h$, a contradiction. Analogously, the sequence of $v$-$t$ bridges contains the arcs of $A$ succeeding $h$ in $p$, since otherwise there would be a fork in the nodes of $p$ succeeding $h$, a contradiction. Therefore any safe sequence is either the bridge extension of some arc in the sequence (or a subsequence of the extension, as dropping arbitrary arcs from a safe sequence yields another safe sequence). From this discussion we can also bound the number of maximal safe sequences.
Corollary 3
There are at most $m$ maximal safe sequences.
In spite of the fact that the number of safe sequences and the number of safe paths are both upper bounded by the number of arcs in the graph, we remark that safe sequences are at least as long as safe paths because any safe path is also a safe sequence by definition.
Bridge extensions have analogous limitations to univocal extensions. Indeed, applying bridge extensions to two distinct arcs may result in the same sequence, and moreover the resulting sequence may not be maximal. Nonetheless, from the discussion above, by extending every arc through its bridges, left and right, we are bound to find all the maximal safe sequences in the graph. We now present a lemma that will simplify the rest of this section. See Appendix 0.A for the proof.
{restatable*}
lemmaextsubstring If $A$ is a non maximal safe sequence computed via bridge extensions then it is a prefix, suffix, or prefix and suffix of a maximal safe sequence.
Theorem 3.1
Given a DAG $G$ with $m$ arcs, there is an $O(m^{2})$-time algorithm computing a set of safe sequences containing all the maximal safe sequences of $G$.
Proof
For every arc $(u,v)\in E(G)$ compute the set of $s$-$u$ and $v$-$t$ bridges. This computation can be done in $O(m)$ time using two DFS executions using, e.g., the algorithm proposed in [CAIRO2021103].
A faster algorithm via dominators.
Another way of framing bridge extensions is via the well known notion of dominators. Given a digraph with a designated start vertex $s$, we say that an arc $(x,y)$ dominates an arc $(z,w)$ if all paths from $s$ to $z$ contain $a$. Clearly, $a$ is an $s$-$z$ bridge if and only if $a$ dominates $b$. There is an extensive study on dominators in the literature, namely that of computing the dominator tree of a general digraph. We refer to [chandak2016survey] for details.
Given a DAG $G=(V,E)$ and an arc $(u,v)\in E$, our goal compute is to compute $s$-$u$ bridges and $v$-$t$ bridges fast. We proceed as follows. First we build the dominator tree $T_{s}$ of $G$ rooted at $s$ and the dominator tree $T_{t}$ of $G$ with all its arcs reversed rooted at $t$. Now, left bridge extensions from an arc $a$ coincide with the traversal from $a\in V(T_{s})$ to $s$ in $T_{s}$. Analogously, the right bridge extensions from $a$ coincide with the traversal from $a\in V(T_{t})$ to $t$ in $T_{t}$. The correctness of the former is clear, and of the former comes from the fact that an arc $b$ is a $v$-$t$ bridge in the original graph if and only if $b$ is a $t$-$v$ bridge in the reversed graph.
A dominators’ related technique that we can use to compute bridge extensions is that of Italiano et al. in [italiano], where they showed the following result. Given a strongly connected directed graph $G$, in $O(m+n)$ time and $O(n)$ auxiliary space it is possible to efficiently answer the following query: given two nodes $u$ and $v$, report all arcs $a\in E(G)$ such that $u$ does not reach $v$ or $v$ does not reach $u$ in $G-a$. Following the terminology of [italiano], we will call such an arc $a$ a separating arc for $u,v$. Importantly, therein they show this can be done in linear time on the number of separating arcs. Then, we can use the techniques of [italiano] together to solve Problem 2.
[…]
Proof
Let $G^{\prime}=(V(G),E(G)\cup(t,s))$ be a strongly connected graph and let $(u,v)\in E(G)$. First we want to show that $a\neq(t,s)$ is a separating arc for $s,u$ in $G^{\prime}$ if and only if it is an $s$-$u$ bridge in $G$, and the argument for $v,t$ separators is symmetric. If $a\neq(t,s)$ is a $s$-$u$ bridge in $G$ then all $s$-$u$ paths in $G$ contain $a$ and thus there is no $s$-$u$ path in $G^{\prime}-a$, as an $s$-$u$ path in $G^{\prime}-a$ would be a $s$-$u$ path in $G$ not containing $a$, contradicting the fact that $a$ is a $s$-$u$ bridge. If $a\neq(t,s)$ is a separating $s,u$ arc then there is no $s$-$u$ path in $G^{\prime}-a$, that is, every $s$-$u$ path in $G^{\prime}$ contains $a$, and thus $a$ is a $s$-$u$ bridge in $G$.
Our algorithm first computes for every arc $(u,v)\in E(G)$ all the arcs that separate $s$ and $u$ in $G^{\prime}$, and those that separate $v$ and $t$ in $G^{\prime}$, thus finding all the $s$-$u$ and $v$-$t$ bridges. Note that the arc $(t,s)$ separates $s,u$ ($v,t$) because all $u$-$s$ ($t$-$v$) paths contain $(t,s)$, so whenever $(t,s)$ is reported as a separating arc we simply do not add it to the bridge sequence.
We analyze separately the time required to compute maximal and non maximal safe sequences. Note that the time spent to compute the separating $s$-$u$ arcs and the separating $v$-$t$ arcs is exactly the length of the safe sequence being computed. Let $s_{1}+s_{2}+\dots+s_{\ell}=o$, where $s_{i}$ denote the length of the $i$-th maximal safe sequence and $\ell\leq m$ by Corollary 3.
For maximal sequences we spend at least $O(o)$ time. In the worst case, each arc of every maximal safe sequence is a core for that sequence, and so we spend at most $s_{1}^{2}+s_{2}^{2}+\dots s_{\ell}^{2}=O(o^{2})$ time to compute all the duplicates of the maximal safe sequences.
By Corollary 3, each extension of an arc that resulted in a non maximal safe sequence is either a prefix, suffix, or prefix and suffix of a maximal safe sequence. In the worst case, every non maximal sequence is a prefix and suffix, as the sum of lengths of all prefixes (or suffixes) of a string of length $k$ is $O(k^{2})$ and that of all strings that are prefixes and suffixes is $O(k^{3})$. Then summing over all safe sequences that are prefixes and suffixes of maximal safe sequences we get $s_{1}^{3}+s_{2}^{3}+\dots+s_{\ell}^{3}=O(o^{3})$.
The drawback of Section 1.0.2 with respect to its analogue Figure 1 for safe paths is that we perform extensions on every arc, instead on only those identifying maximal safe sequences, i.e., the cores. As such, we leave as open problem whether one can compute the maximal sequences in optimal $O(m+o)$ time, where $o$ denotes the length of all maximal safe sequences. We remark that the main ingredient to achieve the optimal running time for the analogous problem for paths (Problem 1) is that we can pinpoint in constant time the cores of the graph by spending only a linear time in a preprocessing step, that of building the contracted graph. A corresponding technique in the context of safe sequences would immediately lead to an optimal time algorithm. Finally, we remark that any technique to build dominator trees fast can be used in our problem, namely the classical results presented in [tarjan1974finding, lengauer1979fast].
4 Experiments
In this section we show experimental results of applying safe paths and safe sequences to scaling two path-finding ILPs, LeastSquares and MinPathError. Due to lack of space, we describe these in detail in Appendix 0.B, together with how we use safe paths and sequences to fix variables in these. We consider these two ILPs as interesting test-cases for our experiments, as they have orthogonal complexity. The former one does not have other variables and constraints apart from the ones encoding solution paths and to linearize some products of variables, see Appendix 0.B; however, it has quadratic terms in its objective function, which generally slows down the solver. The latter one has only linear terms in its objective function, but it has one additional error variable per solution path (making it also more complex than the original ILP for the minimum flow decomposition problem in the error-free case [dias2022fast]).
Choosing the best number $k$ of solution paths in the ILP is a complex issue, and different RNA assembly tools have different approaches, see e.g. the discussion in [cidane]. In this paper we chose to focus on the speed-ups of the safety optimizations, and thus in all ILPs we set $k$ to equal the (arc) width of the DAG, namely the minimum cardinality of a path cover of its arcs. In practice, one can just run the ILPs for different values of $k$, see e.g. [dias2024robust]. Finally, note that any solution to MinPathError is also a path cover, by definition. This is not always the case for LeastSquares, because some arcs may not be covered by any solution path and contribute with the squared weight to the objective function. However, this can be handled also theoretically, see the discussion in Section 5.
We experiment with eight datasets. The first four ones contain splice graphs with erroneous weights created directly from gene annotation [dias2024robust] [^3] Available at https://doi.org/10.5281/zenodo.10775004. These were created starting from the well-known dataset of Shao and Kingsford [shao2017theory], which was also used in several other studies benchmarking the speed of minimum flow decomposition solvers [dias2022fast, kloster2018practical, williams2023subpaths]. The original splice graphs were created in [shao2017theory] from gene annotation from Human, Mouse and Zebrafish. This dataset also contains a set of graphs (SRR020730 Salmon) from human gene annotation, but with flow values coming from a real sequencing sample from the Sequence Read Archive (SRR020730), quantified using the tool Salmon [salmon]. To mimic splice graphs constructed from real read abundances, [dias2024robust] added noise to the flow value of each arc, according to a Poisson distribution. The next four datasets contain graphs constructed by a state-of-the-art tool for RNA transcript discovery from long RNA-seq reads, IsoQuant [isoquant]. Mouse annotated transcript expression was simulated according to a Gamma distribution that resembles real expression profile [isoquant]. From these expressed transcripts, PacBio reads were simulated using IsoSeqSim [isoseqsim] and fed to IsoQuant for graph creation (we call this Mouse PacBio), and ONT reads were simulated using Trans-Nanosim [trans-nanosim], which was modified to perform a more realistic read truncation (see [isoquant]), and fed to IsoQuant for graph creation (we call this Mouse ONT). For the latter one, there is also a simpler version where no read truncation was introduced (Mouse ONT.tr). These are available at [isoquant-graphs]. The last dataset, IsoQuant LRGASP, contains graphs built by IsoQuant from a real read dataset used in the evaluation performed by The Long-read RNA-Seq Genome Annotation Assessment Project Consortium [lrgasp], namely human H1 cell line, ONT cDNA data, id ENCSR016IKV [^4] https://www.encodeproject.org/experiments/ENCSR016IKV/.
The experiments were performed in an AMD 32-core machine with 512GB RAM. The source code of our project is publicly available on Github [^5] https://github.com/algbio/safe-sequences. All our algorithms are implemented in Python3, as are the scripts to produce the experimental results. To solve integer linear programs we use Gurobi’s Python API (version 11) under an academic license. For the first seven datasets we use 4 threads and set up a timeout of 5 minutes in Gurobi’s execution time. Due to the long running of the unoptimized ILPs, for IsoQuant LRGASP we use 32 threads, a timeout of 3 minutes, and we experiment only with the first 5000 graphs out of 17109, and additionally having at most 400 arcs and width at most 60.
Table 1: Speed up metrics with MinPathError.
Width #Solved by all Avg. ILP time (s) Avg. speedup ($\times$)
No safety Safe paths Safe sequences Safe paths Safe sequences
Zebrafish 1-3 15405 0.014 0.010 0.010 1.3 1.3
4-6 239 2.198 0.246 0.111 10.1 13.0
7-9 4 81.099 3.422 0.567 103.1 161.3
10+ 0 - - - - -
Human 1-3 10729 0.022 0.012 0.011 1.7 1.7
4-6 944 3.764 0.298 0.136 14.0 16.8
7-9 66 70.027 3.257 1.742 83.5 103.2
10+ 1 96.390 3.008 1.967 32.0 49.0
Mouse 1-3 12280 0.017 0.010 0.010 1.5 1.5
4-6 749 2.603 0.222 0.140 11.6 13.5
7-9 43 59.249 4.526 4.751 48.1 41.3
10+ 1 295.812 2.691 3.368 109.9 87.8
SRR020730 Salmon 1-3 35069 0.018 0.009 0.009 1.7 1.7
4-6 4496 2.170 0.148 0.131 11.0 11.8
7-9 839 45.547 1.455 1.162 79.4 86.7
10+ 79 68.251 1.718 1.197 113.1 128.7
Mouse PacBio 1-3 14256 0.014 0.011 0.011 1.4 1.4
4-6 1376 0.239 0.081 0.046 4.3 4.8
7-9 181 4.812 0.185 0.170 29.3 29.5
10+ 52 20.143 0.731 0.719 68.1 78.5
Mouse ONT 1-3 18527 0.015 0.010 0.009 1.6 1.6
4-6 3083 0.203 0.102 0.039 4.3 4.6
7-9 755 2.382 0.202 0.135 14.8 16.8
10+ 370 18.003 0.485 0.353 82.1 83.7
Mouse ONT.tr 1-3 18029 0.012 0.009 0.009 1.4 1.5
4-6 3028 0.187 0.055 0.038 3.2 3.5
7-9 796 2.332 0.263 0.165 11.4 13.9
10+ 400 15.213 0.374 0.302 56.8 62.4
LRGASP 1-10 2919 0.070 0.013 0.013 2.0 2.0
11-20 457 12.610 0.353 0.365 62.3 66.2
21-30 169 48.853 1.446 1.859 112.6 120.6
31-45 44 84.038 1.033 1.211 155.3 161.4
46-60 5 91.767 0.838 0.614 113.4 153.1
  • In each dataset we bin the graphs based on their width, and in this table we keep the graphs that were solved by all ILP versions; column “#Solved by all” indicates the number of such graphs in each bin. Inside each bin, we show the average (i.e. mean) of the metric of the graphs in the bin. The speed-ups are computed as follows: for each graph, we divide the time the original ILP takes (with no safety information) by the time the safety optimized ILP takes (with paths, and sequences, respectively). Then, inside a bin, we show the average of these speed-ups. Note that this value is different from the one obtained by just dividing the average times in each bin. For the first seven datasets we use a timeout of 5 minutes and 4 threads, and for IsoQuant LRGASP we use a timeout of 3 minutes and 32 threads.
In Table 1 we show the speed-ups of the safety-optimized ILPs for MinPathError with respect to the original one. When the width of the graph is small, the speed-ups are also small; however, such graphs are fast to solve in the first place. The challenging cases seem to be the graphs with larger widths, as we observe that the ILP solver with no safety performs poorly and it becomes slower as the width increases. The average speed-up of the solver significantly increases as the width of the graph increases, becoming more than 150$\times$ on IsoQuant LRGASP. Similar speed-ups are achieved also for LeastSquares (see Table 3 in Appendix 0.C), with some speed-ups of more than 300$\times$ (and up to 1000$\times$) on the mouse datasets of IsoQuant. Moreover, these speed-up are obtained almost for free, in the sense that the time needed to compute safe paths and safe sequences is negligible for all datasets (i.e. below 0.06 seconds, and often much smaller) independently of the width (see Table 2, as well as Table 4 in Appendix 0.C).
While Tables 1 and 3 focus on the graphs solved by all three ILP variants, in Tables 2 and 4 we show the number of graphs solved by each ILP, and the average time the ILP takes on the graphs that are solved. We observe that the safety-optimized ILPs solve more graphs than the original ILPs. For example, for MinPathError, on SRR020730 Salmon, when the width is at least 10, we solve at least 3.5 times more graphs, while being at least 3 times faster. More significantly, for MinPathError on IsoQuant LRGASP, when the width is larger than 46, we solve 34 times more graphs, while being about 9 times faster.
In Tables 2 and 4 we also quantify the amount of safety information that we use to optimize the ILPs. To have a uniform metric for both paths and sequences, we measured this as the percentage of variables that we fix with safe paths and safe sequences, respectively, out of all the $m\cdot k$ binary variables used by the ILP to encode the $k$ solution paths (as in [acceleratingILP]). For the interesting graphs with larger width, we can fix around 10% of the binary variables of the ILP. We find it remarkable that we can obtain the significant speed-ups discussed above with such relatively little use of safety information. Moreover, the amount of fixed variables depends only on the graph structure, not on the path-finding problem formulation. Thus, the percentage of fixed variables indicates the reduction in search space for solution paths for any path-finding ILP model. We can thus speculate that significant speed-ups can be obtained also for more complex ILPs.
Finally, when comparing safe paths and safe sequences on MinPathError, we generally see slight advantages of sequences with respect to paths (with few exceptions). In terms of average ILP time, the notable improvements are in the “medium” width ranges. Overall, the former one also solves at least as many instances in each dataset. On LeastSquares the advantages of safe sequences are even more pronounced.
Table 2: Metrics on safety-preprocessing time, solved instances, and fixed variables, with MinPathError.
Width #Graphs Avg. preproc time (s)) #Solved (Avg. time (s)) Avg. fixed vars (%)
Safe paths Safe sequences No safety Safe paths Safe sequences Safe paths Safe sequences
Zebrafish 1-3 15418 0.003 0.004 15405 (0.014) 15405 (0.013) 15405 (0.013) 85.4 87.5
4-6 239 0.006 0.010 239 (2.198) 239 (0.252) 239 (0.121) 22.3 29.5
7-9 6 0.013 0.028 4 (81.099) 6 (3.299) 6 (0.780) 11.7 15.3
10+ 1 0.015 0.028 0 (-) 1 (3.582) 1 (3.681) 8.6 13.1
Human 1-3 10732 0.003 0.004 10729 (0.022) 10729 (0.014) 10729 (0.015) 75.9 78.9
4-6 947 0.006 0.009 944 (3.764) 947 (0.307) 947 (0.149) 18.5 24.7
7-9 92 0.010 0.015 66 (70.027) 91 (5.391) 92 (3.925) 10.4 14.1
10+ 12 0.014 0.024 1 (96.390) 10 (63.760) 9 (27.527) 9.1 11.7
Mouse 1-3 12304 0.003 0.004 12280 (0.017) 12280 (0.013) 12280 (0.014) 82.0 84.4
4-6 749 0.006 0.010 749 (2.603) 749 (0.229) 749 (0.149) 20.3 27.4
7-9 60 0.010 0.021 43 (59.249) 58 (5.273) 60 (5.578) 11.3 17.5
10+ 9 0.012 0.022 1 (295.812) 6 (13.283) 7 (36.930) 9.9 15.2
SRR020730 Salmon 1-3 35069 0.002 0.002 35069 (0.018) 35069 (0.010) 35069 (0.011) 82.9 84.7
4-6 4497 0.006 0.009 4496 (2.170) 4497 (0.155) 4497 (0.141) 16.1 20.4
7-9 1008 0.009 0.014 839 (45.547) 1008 (2.476) 1007 (1.862) 9.2 11.9
10+ 296 0.014 0.022 79 (68.251) 284 (21.334) 287 (22.698) 6.1 7.9
Mouse PacBio 1-3 14256 0.003 0.004 14256 (0.014) 14256 (0.014) 14256 (0.015) 82.9 83.5
4-6 1376 0.007 0.009 1376 (0.239) 1376 (0.088) 1376 (0.055) 30.7 33.3
7-9 182 0.010 0.013 181 (4.812) 182 (0.202) 182 (0.184) 18.2 20.5
10+ 63 0.029 0.036 52 (20.143) 62 (3.310) 62 (3.008) 12.1 13.3
Mouse ONT 1-3 18527 0.002 0.003 18527 (0.015) 18527 (0.012) 18527 (0.012) 81.0 81.4
4-6 3083 0.006 0.009 3083 (0.203) 3083 (0.107) 3083 (0.048) 28.6 30.2
7-9 762 0.009 0.013 755 (2.382) 761 (0.326) 762 (0.195) 16.9 18.1
10+ 442 0.020 0.024 373 (18.680) 437 (0.728) 437 (0.622) 10.2 11.1
Mouse ONT.tr 1-3 18029 0.002 0.003 18029 (0.012) 18029 (0.012) 18029 (0.012) 80.6 81.1
4-6 3028 0.006 0.009 3028 (0.187) 3028 (0.061) 3028 (0.047) 28.6 30.6
7-9 803 0.010 0.015 796 (2.332) 803 (0.428) 803 (0.201) 17.3 19.0
10+ 476 0.021 0.027 401 (15.180) 472 (1.034) 472 (0.878) 10.5 11.5
LRGASP 1-10 2919 0.002 0.002 2919 (0.070) 2919 (0.015) 2919 (0.015) 68.9 69.1
11-20 492 0.011 0.014 457 (12.610) 490 (0.646) 490 (0.419) 7.3 7.6
21-30 310 0.019 0.025 170 (48.978) 299 (2.955) 298 (3.276) 3.9 4.1
31-45 293 0.029 0.038 44 (84.038) 272 (8.007) 275 (8.229) 2.6 2.7
46-60 214 0.051 0.059 5 (91.767) 171 (10.419) 176 (11.674) 1.9 2.0
  • In each dataset, graphs are binned by width. Column “#Graphs” shows the total number of input graphs inside each bin. In all the other metrics, we show averages with respect to the graphs that were solved by the respective ILP. Column “Avg. preproc time” shows the average times needed to compute safe paths, and safe sequences, respectively. Column “#Solved (Avg. time)” shows the number of graphs that were solved by each ILP, and the average time taken for all solved graphs. Column “Avg. fixed vars” shows the percentage of variables set to 1 via safe paths and safe sequences, respectively, out of all the $m\cdot k$ binary variables of the ILP. For the first seven datasets we use a timeout of 5 minutes and 4 threads, and for IsoQuant LRGASP we use a timeout of 3 minutes and 32 threads.
5 Discussion and conclusions
Despite the many applications of path-finding ILPs in RNA transcript assembly problems, most optimizations introduced so far have focused on the error-free case. In this paper we showed for the first time that speed-ups of up to two orders of magnitude are possible also for DAGs with arbitrary erroneous weights.
To obtain these, we used the recent introduction of polynomially-sized ILPs for path-finding problems from [dias2022fast, jumper], and the approach from [acceleratingILP] of reducing the search space of these ILPs using the safety information of the graph. However, as opposed to [acceleratingILP], our results apply to any path-finding ILP, as long as its solution paths are a path cover of the DAG. To this purpose, we gave a linear-time algorithm computing all safe paths for paths covers, generalizing the well-known Y-to-V reduction. Moreover, we also considered safety information in the form of safe sequences for path covers, inspired by [and_or_quant] which studied a similar concept (but again for the error-free case of flow decompositions).
Our safety optimizations might be incorrect if there are arcs erroneously present in the graph (e.g. with low abundance), in which case an ILP formulation might choose to avoid these arcs and thus admit an optimal solution that is not a path cover. As a heuristic, one could compute safe paths and sequences only for path covers of a graph where these arcs are removed. As a more principled approach, our methodology could be generalized using the notion of subset covering [hydrostructure]. Instead of considering path covers of all arcs, we consider path covers of a subset of arcs that we trust to correctly appear in the graph. We believe that our theoretical results can be generalized to this case, as was done for more complex notions of covers in [hydrostructure], and we leave this for future work.
In this paper we considered the well-known LeastSquares model and the more recent but more accurate MinPathError model. To precisely measure the speed-ups that safe paths and safe sequences allow, we deliberately did not include other constraints in these models; instead, we focused on testing these optimization on a variety of datasets. While there are many other models that are relevant in practice (see e.g. [dias2022fast, dias2024robust, tomescu2015explaining, jumper, scallop2] and the references cited above), our speed-ups are obtained solely from the graph structure, without any other knowledge of the problem formulation, nor of the graph weights. In other words, the simplification and reduction in the ILP search space that safety carries over to any other path-finding ILP. As such, we hope that our optimization can be employed at the core of future ILP-based RNA transcript assembly tools that scale to large and complex graphs.
Acknowledgements.
We are extremely grateful to Andrey Prjibelski for producing the IsoQuant graphs and to Manuel Cáceres for fruitful discussions on safe sequences. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 851093, SAFEBIO), and from the Research Council of Finland grants No. 346968, 358744.
[…]
Appendix 0.A Missing proofs
[…]
Proof
We can assume that $k\geq 2$ because in an $s$-$t$ DAG every $s$-$t$ path contains at least two nodes.
()($\Rightarrow$)( ⇒ ) Suppose by contradiction that $p$ is a safe path and has an inner node $v_{j}$ with $d^{+}_{v_{j}}\geq 2$ not preceding an inner node $v_{i}$ with $d^{-}_{v_{i}}\geq 2$, $1<i\leq j<k$, and let $P$ be a path cover. We will show that $P$ can be modified so that none of its paths will contain $p$ as subpath, which will contradict our assumption that $p$ is safe. Let $a_{i}=u_{i}v_{i}$ and $a_{j}=v_{j}u_{j}$ be arcs where $u_{i},u_{j}\notin p$ and consider an $s$-$t$ path $q$ that contains $a_{i}$ and $a_{j}$. Note that $q$ does not contain $p$ as a subpath, otherwise there would be a cycle. Now, for any path $r\in P$ containing $p$ as a subpath we can splice $q$ with $r$ by building $q_{1}=q\bar{u_{i}}\bar{v_{i}}r$ and $q_{2}=r\bar{v_{j}}\bar{u_{j}}q$. Since $p$ is not a subpath of $q_{1},q_{2}$ and $E(r)\subseteq E(q_{1})\cup E(q_{2})$, we can update $P$ by removing $r$ and adding $q_{1},q_{2}$ while maintaining $P$ a path cover. Doing this for every path in $P$ that contains $p$ as a subpath shows that $p$ is unsafe, a contradiction.
()($\Leftarrow$)( ⇐ ) Suppose by contradiction that $p$ is unsafe but any inner node with outdegree at least two precedes any inner node with indegree at least two. Let $v_{j}$ be the first inner node of $p$ with indegree at least two, or $v_{k}$ if $p$ has no such inner node, and let $v_{i}$ be its last inner node with oudegree at least two, or $v_{1}$ if $p$ has no such inner node. Then there is an arc $h=v_{i^{\prime}}v_{j^{\prime}}$ in $p$ such that $i\leq i^{\prime}<j^{\prime}\leq j$. Now, any $s$-$t$ path covering $h$ has the form $q_{1}v_{i^{\prime}}v_{j^{\prime}}q_{2}$ where $q_{1}$ is an $s$-$v_{i^{\prime}}$ path and $q_{2}$ is a $v_{j^{\prime}}$-$t$ path. In particular, $q_{2}$ contains the suffix of $p$ succeeding $v_{i^{\prime}}$, since every inner node of $p$ following $v_{i^{\prime}}$ has outdegree exactly one. Symmetrically, $q_{1}$ contains the prefix of $p$ preceding $v_{j^{\prime}}$. since every inner node of $p$ preceding $v_{j^{\prime}}$ has indegree exactly one. Therefore, any $s$-$t$ path covering $h$ must contain $p$ as a subpath and thus $p$ is safe, a contradiction.
Figure 2: The green path $p=v_{1}\dots v_{k}$ is unsafe, since $v_{j}$ does not precede $v_{i}$ and $d^{+}_{v_{j}}\geq 2$ and $d^{-}_{v_{i}}\geq 2$. Any $s$-$t$ path of a path cover containing $p$ (e.g., the green path) can be replaced by the two orange paths. Following the proof of Section 2, the bottom orange path corresponds to path $q_{1}$ and the top orange path corresponds to path $q_{2}$.
[…]
Proof
()($\Rightarrow$)( ⇒ ) Suppose by contradiction that $A$ is safe but any $\textsc{h}(a_{1})$-$\textsc{t}(a_{\ell})$ path $p$ containing all arcs in $a_{2},\dots,a_{\ell-1}$ has a fork $v_{j}$ not preceding a merge $v_{i}$. Let $P$ be an $s$-$t$ path cover and let $r=v_{1}\dots v_{k}$ be any path in $P$ that covers $A$. We will modify every such path $r$ so that $A$ is no longer a subsequence of $r$ while maintaining $P$ a path cover.
Note that $r=s\dots\textsc{t}(a_{1})\,p\,\textsc{h}(a_{\ell})\dots t$ where $p=\textsc{h}(a_{1})\dots v_{i}p^{\prime}v_{j}\dots\textsc{t}(a_{\ell})$ and $V(p)\subseteq R_{A}$. Possibly $v_{i}=v_{j}$, and in that case $p^{\prime}$ is the empty path and the concatenation $v_{i}p^{\prime}v_{j}$ results in $v_{i}$. Further, since $v_{j}$ does not precede $v_{i}$ in $p$, we have that $v_{i}\in R_{a_{i^{\prime}}a_{{i^{\prime}}+1}}$ and $v_{j}\in R_{a_{j^{\prime}}a_{{j^{\prime}}+1}}$ for $11\leq i^{\prime}\leq j^{\prime}\leq\ell-$ 1, and so there are nodes $u_{i}\notin R_{a_{i^{\prime}}a_{{i^{\prime}}+1}}$, $u_{j}\notin R_{a_{j^{\prime}}a_{{j^{\prime}}+1}}$ and paths $p_{i}=u_{i}\dots v_{i}$, $p_{j}=v_{j}\dots u_{j}$ witnessing that $v_{i},v_{j}$ are merge and fork nodes, respectively. Observe that possibly $u_{i},u_{j}\in R_{A}$. Let $q=s\dots p_{i}p^{\prime}p_{j}\dots t$ and build the paths $q_{1}=q\bar{u_{i}}p_{i}\bar{v_{i}}r$ and $q_{2}=r\bar{v_{j}}p_{j}\bar{u_{j}}q$. Since $A$ is not a subsequence of $q_{1},q_{2}$ and $E(r)\subseteq E(q_{1})\cup E(q_{2})$, we can modify $P$ by removing $r$ and adding $q_{1},q_{2}$ while maintaining $P$ a path cover. Doing this splicing operation for every path that covers $A$ contradicts the initial assumption that $A$ is safe.
()($\Leftarrow$)( ⇐ ) Suppose by contradiction that $A$ is unsafe but there is a $\textsc{h}(a_{1})$-$\textsc{t}(a_{\ell})$ path $p=v_{1}\dots v_{k}$ containing the sequence $a_{2},\dots,a_{\ell-1}$ where any fork of $A$ in $p$ precedes any merge of $A$ in $p$. Let $v_{j}$ be the first merge of $p$, or $v_{k}$ if $p$ has no merge, and let $v_{i}$ be its last fork, or $v_{1}$ if $p$ has no fork. Then there is an arc $h=v_{i^{\prime}}v_{j^{\prime}}$ in $p$ such that $i\leq i^{\prime}<j^{\prime}\leq j$ (note that possibly $h\in A$). Now, $h$ must be covered by some $s$-$t$ path $q$ of a path cover, and any such path must contain every arc of $A$ that succeeds $h$ in $p$, since otherwise there would exist a fork in $p$ succeeding $v_{i^{\prime}}$, but by the choice of $v_{i}$ this cannot happen. In other words, all arcs of $A$ succeeding $h$ in $p$ are $v_{j^{\prime}}$-$t$ bridges. Likewise, $q$ must contain every arc of $A$ that precedes $h$ in $p$, and those arcs are $s$-$v_{i^{\prime}}$ bridges. To conclude, any $s$-$t$ path containing $h$ contains $A$, and so $A$ is safe, a contradiction.
Figure 3: The green sequence $A=a_{1}\dots a_{5}$ is unsafe, as there is a merge in $R_{a_{2}a_{3}}$ and a fork in $R_{a_{3}a_{4}}$. Any $s$-$t$ path of a path cover containing $A$ can be replaced by the two orange paths. Following the proof of Section 3, the bottom orange path corresponds to path $q_{1}$ and the top orange path corresponds to path $q_{2}$. Note that the subgraphs $R_{a_{i}a_{i+1}}$ for $i\in\{1,\dots,4\}$ can be arbitrarily complex.
(a) A configuration in a graph $G$ with a node $v$ having a unique in-neighbor $u$, and out-neighbors $w_{1}$ and $w_{2}$.
(b) The result of applying the Y-to-V reduction on $v$: node $v$ is removed, and arcs are added from $u$ to both $w_{1}$ and $w_{2}$, with the resulting graph being $G^{\prime}$.
Figure 4: An example showing that the Y-to-V reduction misses safe paths for path covers of a DAG. On the left, in $G$ there are four safe paths, shown in green. On the right, after applying the Y-to-V reduction, in $G^{\prime}$ we still have four safe paths made up of a single arc, also shown in green. If we keep track that the newly added arcs $(u,w_{1})$ and $(u,w_{2})$ in $G^{\prime}$ correspond to the original paths $uvw_{1}$ and $uvw_{2}$ in $G$, then the two safe paths starting from $u$ in $G^{\prime}$ can be expanded into the two original safe paths $uvw_{1}$ and $uvw_{2}$ in $G$. However, this does not hold for the two safe paths ending at $u$ in $G^{\prime}$, because the arcs entering $u$ are unchanged by the Y-to-V reduction. Thus, these two safe path in $G^{\prime}$, when mapped to the original graph $G$, miss the arc $(u,v)$.
[…]
Proof
Let $B$ be a maximal safe sequence containing $A$. Since $A$ is computed via bridge extensions there is an arc $h_{A}\in A$ whose bridge extension is $A$, and since $B$ is maximally safe there is an arc $h_{B}\in B$ whose bridge extension is $B$. If $h_{A}=h_{B}$ then they have the same bridge extension, so $h_{A}\neq h_{B}$.
Using Section 3, let $p$ be a path where every fork of $B$ precedes every merge of $B$ in $p$. Since $A$ is a subsequence of $B$, a merge (fork) node in $R_{b_{i}b_{i+1}}$ (for any $i$) with respect to $A$ is also a merge (fork) node with respect to $B$, and thus every fork of $A$ in $p$ precedes every merge of $A$ in $p$ (note that $B$ has at least two arcs, so $R_{b_{i}b_{i+1}}$ is well defined). If $h_{B}$ precedes $h_{A}$ in $p$ then, by construction of bridge extensions, the right bridge extension of $h_{A}$ coincides with the right bridge extension of $B$ restricted to the arcs succeeding $h_{A}$, i.e., there are no gaps in this suffix of $A$ with respect to $B$. Moreover, if between $h_{B}$ and $h_{A}$ there is no merge node with respect to $A$ then the left bridge extension of $h_{A}$ spans over the arcs between $h_{B}$ and $h_{A}$, and thus the extension would continue through $h_{B}$, implying that they would have the same left bridge extension. Then there is a merge node between $h_{B}$ and $h_{A}$, and so let $v\in p$ be the rightmost such node. Again, by construction of bridge extensions, the left bridge extension of $h_{A}$ contains every arc of $B$ between $v$ and $\textsc{t}(h_{A})$ and stops at $v$. If there is an $s$-$v$ with no internal node in common with $p\setminus\{s\}$, we have that the subsequence between $v$ and $\textsc{t}(h_{A})$ concatenated with $h_{A}$ and this concatenated with the right bridge extension of $A$ results in $A$, and this in turn is a suffix of $B$. If the path witnessing the fact that $v$ is a merge intersects $p\setminus\{s\}$, then by the definition of bridge extension every arc of $B$ preceding the first node of this path will also be in $A$, and thus $A$ is a prefix and suffix of $B$. If $h_{A}$ precedes $h_{B}$ we can argue analogously to conclude that $A$ is a prefix or a prefix and suffix of $B$.
Appendix 0.B Scaling path-finding ILPs using safety
0.B.0.1 Polynomially-size path-finding ILPs.
In this section we describe our application of safe paths and sequences to simplifying (and consequently, speeding up) path-finding ILPs. As already mentioned, Dias et al. [dias2022fast], and Sashittal et al. [jumper] showed that such type of problems admit polynomially-sized ILPs, improving on previous ILP formulations [cidane, class2, cliiq, isoinfer, isolasso, multitrans, nsmap, ssp, translig] which are potentially exponentially-sized. These work in three parts: (i) $k$ paths are encoded via suitable variables and constraints; (ii) additional constraints (and variables) are added to encode when these paths form a solution; (iii) if necessary, these variables are then used in the formulation of the objective function of the ILP (e.g. minimization or maximization). Such an ILP can also have a wrapper in which one searches for different numbers $k$ of paths, for each $k$ solves the above-mentioned ILP, and then chooses that value of $k$ optimizing some quantity. In what follows, we review two specific ILPs, and we refer the reader to the original papers [dias2022fast, jumper] for further details.
For part (i), for each solution path $i$ that needs to be encoded ($i\in\{1,\dots,k\}$), both [dias2022fast, jumper] introduce a binary variable $x_{uvi}$ associated to each arc $(u,v)$ of the DAG, with the interpretation that $x_{uvi}$ equals 1 if and only if the arc $(u,v)$ belongs to the $i$-th solution path. Since the input graph is DAG, this can be guaranteed by stating that the sum of all $x_{sui}$ variables on arcs $(x,u)$ exiting the source $s$ is exactly 1, and for all nodes $v$ (different from $s$ and $t$) the sum of the binary variables $x_{uvi}$ of arcs $(u,v)$ entering $v$ is equal to the sum of the binary variables $x_{vwi}$ of arcs $(v,w)$ exiting $v$, see [dias2022fast].
0.B.0.2 Fixing ILP variables via safe paths and sequences.
In this paper, we apply safe paths and sequences to fix some $x_{uvi}$ variables to $1$, thus reducing the search space of the ILP solver, using the approach of Grigorjew et al. [acceleratingILP]. Let $P_{1},\dots,P_{t}$ be paths in the input DAG such that: (a) each $P_{i}$, $i\in\{1,\dots,t\}$, must appear in some solution path of the ILP, and (b) there exists no path in the DAG containing two distinct $P_{i}$ and $P_{j}$. Grigorjew et al. [acceleratingILP] noticed that if these two properties hold, then $P_{1},\dots,P_{t}$ must appear in $t$ distinct paths of among the $k$ solution paths of the ILP. As such, one can, without loss of generality, assign each $P_{i}$ to the $i$-th solution path. That is, for each $i\in\{1,\dots,t\}$, we set $x_{uvi}=1$ for each arc $(u,v)$ of $P_{i}$.
When using sequences to set binary variables, we can proceed in a completely analogous manner. Assume $S_{1},\dots,S_{t}$ are sequences tha must appear in some solution path, and such that no path in the DAG contains distinct $S_{i}$ and $S_{j}$. As such, we can correctly assign $S_{i}$ to the $i$-th solution path of the ILP. Namely, we can analogously set $x_{uvi}=1$ for each arc $(u,v)$ of $S_{i}$.
Finally, if it holds that the ILP solution paths cover all arcs (i.e. they are a path cover of the arcs), then we can use safe paths (or safe sequences, respectively) in the above procedure to ensure that the ILP still has the same set of optimal solutions. To find suitable sets of paths and sequences, and to have a large number of binary variables set to 1 in this manner, we follow the approach of [acceleratingILP]. Namely, for every arc $(u,v)$, we record the longest safe path (safe sequence, respectively) containing $(u,v)$, and use its length as the weight of $(u,v)$. Then, we find a maximum-weight antichain of arcs in the DAG, namely a set of arcs $e_{i},\dots,e_{t}$ such that from no $e_{i}$ we can reach some other $e_{j}$, $i,j\in\{1,\dots,t\}$, and its total weight is maximum among all antichains, using the min-flow reduction of [rival2012graphs]. Then, for each $e_{i}$ in the maximum-weight antichain, we take the recorded longest safe path $P_{i}$ (or longest safe sequence $S_{i}$, respectively). To see that there can be no path in the DAG containing distinct $P_{i}$ and $P_{j}$ (or $S_{i}$ and $S_{j}$, respectively), notice that such path would contain both the corresponding arcs $e_{i}$ and $e_{j}$, which contradicts the fact that $e_{i}$ and $e_{j}$ belong to an antichain.
For the computation of safe paths we implemented the $O(mn)$-time algorithm described in Section 2. For the computation of safe sequences we implemented the $O(m^{2})$-time algorithm described in Section 3. Although they are not the most optimal algorithms, they serve the purpose of our experiments. Moreover, we experimentally observed that the safety overhead introduced by the execution of these algorithms is negligible.
0.B.0.3 Two path-finding ILPs.
In this section we describe two path-finding ILP that we chose to experiment with. For both, we assume that for every arc $(u,v)$ of the graph we have an associated weight $w(u,v)$. We also assume part (i) discussed above, namely, the $x_{uvi}$ binary variables, and the constraints discussed above to ensure they induce paths. The overall goal is to find the “best” $k$ such paths, and their associated weights $w_{1},\dots,w_{k}$.
We selected a classical least-squares model (which we call LeastSquares) that is at the core of several RNA assembly tools e.g. [isolasso, cidane, ryuto, slide, ireckon, traph]. This minimizes the sum of squared differences between the weight of an arc and the weights of the solution paths passing through the arc. The ILP formulation requires no additional constraints, and the ILP objective function is: [^6] Note that the terms $x_{uvi}w_{i}$ are not linear, but can be linearized using a simple technique, see [dias2022fast, dias2024robust].
$\min\sum_{(u,v)\in E}\left(w(u,v)-\sum_{i=1}^{k}x_{uvi}w_{i}\right)^{2}.$ (1)
While this model is popular, it has quadratic terms in its objective function, which makes it harder to solve. Therefore, as second model (which we call MinPathError), we chose one that was recently introduced in [dias2024robust] and shown to be more accurate than LeastSquares on the datasets considered therein. In this model, errors are accounted not at the level of individual arcs, but at the level of paths. Namely, for every solution path $P_{i}$ we have an associated error (or slack) $\rho_{i}$ which intuitively corresponds to the maximum change in coverage across it. Then, for $k$ paths to be a solution to the ILP, we require that the absolute difference between the weight of an arc and the weights of the solution paths passing through the arc must be below the sum of the errors (or slacks) of the paths passing through the arc. Formally, we add the constraint: [^7] Note again that the terms $x_{uvi}w_{i}$ and $x_{uvi}\rho_{i}$ are not linear, but can be linearized using a simple technique, see [dias2024robust].
[…]
Appendix 0.C Experimental results with the Least-Squares model
Table 3: Speed up metrics with LeastSquares.
Width #Solved by all Avg. ILP time (s) Avg. speedup ($\times$)
No safety Safe paths Safe sequences Safe paths Safe sequences
Zebrafish 1-3 15405 0.029 0.014 0.013 1.6 1.7
4-6 224 12.983 0.324 0.249 38.5 52.6
7-9 0 - - - - -
10+ 0 - - - - -
Human 1-3 10729 0.042 0.018 0.016 2.0 2.2
4-6 865 17.418 0.418 0.329 38.2 43.3
7-9 3 89.071 1.725 4.763 199.8 169.5
10+ 0 - - - - -
Mouse 1-3 12280 0.033 0.015 0.014 1.8 1.9
4-6 698 15.287 0.644 0.399 38.0 44.4
7-9 0 - - - - -
10+ 0 - - - - -
SRR020730 Salmon 1-3 35069 0.028 0.015 0.014 1.8 1.9
4-6 4420 13.691 0.397 0.359 26.4 29.1
7-9 255 126.883 1.057 1.022 213.3 244.7
10+ 0 - - - - -
Mouse PacBio 1-3 14256 0.024 0.008 0.008 2.3 2.4
4-6 1357 2.166 0.081 0.044 36.7 41.4
7-9 150 32.551 0.320 0.109 553.8 581.4
10+ 18 106.511 0.169 0.095 1189.7 1398.8
Mouse ONT 1-3 18527 0.013 0.007 0.007 1.5 1.5
4-6 3079 1.205 0.050 0.036 19.9 20.6
7-9 706 15.583 0.214 0.089 269.7 312.8
10+ 228 34.496 0.434 0.244 385.3 426.7
Mouse ONT.tr 1-3 18029 0.013 0.007 0.007 1.4 1.4
4-6 3017 1.639 0.089 0.033 23.3 25.8
7-9 733 12.798 0.456 0.095 181.3 194.7
10+ 255 30.675 0.177 0.169 318.0 328.7
LRGASP 1-10 2908 0.570 0.023 0.017 6.3 7.7
11-20 184 20.342 0.872 0.805 127.3 155.5
21-30 33 34.803 10.054 11.853 103.4 115.9
31-45 25 38.809 22.294 21.558 12.4 12.1
46-60 16 52.585 19.979 20.637 12.2 13.3
  • In each dataset we bin the graphs based on their width, and in this table we keep the graphs that were solved by all ILP versions; column “#Solved by all” indicates the number of such graphs in each bin. Inside each bin, we show the average (i.e. mean) of the metric of the graphs in the bin. The speed-ups are computed as follows: for each graph, we divide the time the original ILP takes (with no safety information) by the time the safety optimized ILP takes (with paths, and sequences, respectively). Then, inside a bin, we show the average of these speed-ups. Note that this value is different from the one obtained by just dividing the average times in each bin. For the first seven datasets we use a timeout of 5 minutes and 4 threads, and for IsoQuant LRGASP we use a timeout of 3 minutes and 32 threads.
Table 4: Metrics on solved instances, fixed variables, and safety-preprocessing time with LeastSquares.
Width #Graphs Avg. preproc time (s)) #Solved (Avg. time (s)) Avg. fixed vars (%)
Safe paths Safe sequences No safety Safe paths Safe sequences Safe paths Safe sequences
Zebrafish 1-3 15405 0.003 0.004 15405 (0.029) 15405 (0.017) 15405 (0.017) 85.4 87.5
4-6 239 0.006 0.010 224 (12.983) 239 (1.311) 239 (0.503) 22.3 29.5
7-9 6 0.013 0.028 0 (-) 6 (38.952) 6 (5.119) 11.7 15.3
10+ 1 - 0.028 0 (-) 0 (-) 1 (24.383) - 13.1
Human 1-3 10729 0.003 0.004 10729 (0.042) 10729 (0.021) 10729 (0.020) 75.9 78.9
4-6 947 0.006 0.009 865 (17.418) 945 (1.043) 947 (0.593) 18.5 24.7
7-9 92 0.009 0.015 4 (67.110) 80 (20.213) 84 (18.095) 10.8 14.4
10+ 12 0.012 0.022 0 (-) 4 (151.674) 5 (156.942) 9.3 11.3
Mouse 1-3 12280 0.003 0.004 12280 (0.033) 12280 (0.018) 12280 (0.018) 82.0 84.4
4-6 749 0.006 0.010 698 (15.287) 749 (0.948) 749 (0.674) 20.3 27.4
7-9 60 0.010 0.019 0 (-) 47 (31.240) 48 (30.097) 11.6 16.2
10+ 9 0.011 0.018 0 (-) 2 (38.270) 4 (42.273) 9.7 14.7
SRR020730 Salmon 1-3 35069 0.002 0.002 35069 (0.028) 35069 (0.017) 35069 (0.016) 82.9 84.7
4-6 4497 0.006 0.009 4420 (13.691) 4497 (0.533) 4497 (0.487) 16.1 20.4
7-9 1008 0.009 0.016 255 (126.883) 997 (11.624) 996 (10.272) 9.2 11.9
10+ 296 0.014 0.025 1 (7.300) 199 (39.681) 201 (45.535) 6.4 8.3
Mouse PacBio 1-3 14256 0.002 0.003 14256 (0.024) 14256 (0.011) 14256 (0.012) 82.9 83.5
4-6 1376 0.005 0.007 1357 (2.166) 1375 (0.129) 1376 (0.075) 30.7 33.3
7-9 182 0.009 0.012 150 (32.551) 182 (2.340) 182 (1.193) 18.2 20.5
10+ 63 0.012 0.017 18 (106.511) 56 (0.712) 56 (0.473) 13.0 14.2
Mouse ONT 1-3 18527 0.002 0.003 18527 (0.013) 18527 (0.009) 18527 (0.010) 81.0 81.4
4-6 3083 0.005 0.007 3079 (1.205) 3083 (0.056) 3083 (0.044) 28.6 30.2
7-9 762 0.008 0.010 706 (15.583) 759 (0.351) 760 (0.210) 16.9 18.1
10+ 442 0.015 0.019 231 (35.117) 431 (1.390) 433 (0.919) 10.3 11.1
Mouse ONT.tr 1-3 18029 0.002 0.003 18029 (0.013) 18029 (0.009) 18029 (0.010) 80.6 81.1
4-6 3028 0.005 0.007 3017 (1.639) 3027 (0.135) 3028 (0.042) 28.6 30.6
7-9 803 0.009 0.012 734 (12.858) 801 (1.335) 800 (0.274) 17.3 19.0
10+ 476 0.017 0.021 259 (30.952) 458 (2.257) 459 (1.737) 10.6 11.6
LRGASP 1-10 2919 0.002 0.003 2908 (0.570) 2919 (0.027) 2919 (0.022) 68.9 69.1
11-20 492 0.013 0.019 185 (20.278) 476 (4.258) 477 (3.572) 7.4 7.7
21-30 310 0.022 0.029 43 (36.550) 244 (11.168) 244 (10.267) 4.1 4.4
31-45 293 0.033 0.043 54 (40.339) 162 (14.458) 166 (15.692) 2.8 2.9
46-60 214 0.053 0.070 53 (64.062) 87 (24.620) 90 (27.766) 2.0 2.1
  • In each dataset, graphs are binned by width. Column “#Graphs” shows the total number of input graphs inside each bin. In all the other metrics, we show averages with respect to the graphs that were solved by the respective ILP. Column “Avg. preproc time” shows the average times needed to compute safe paths, and safe sequences, respectively. Column “#Solved (Avg. time)” shows the number of graphs that were solved by each ILP, and the average time taken for all solved graphs. Column “Avg. fixed vars” shows the percentage of variables set to 1 via safe paths and safe sequences, respectively, out of all the $m\cdot k$ binary variables of the ILP. For the first seven datasets we use a timeout of 5 minutes and 4 threads, and for IsoQuant LRGASP we use a timeout of 3 minutes and 32 threads.